OpenStudy (anonymous):

How many grams of methane gas (CH4) need to be combusted to produce 18.2 L water vapor at 1.2 atm and 275 K? Show all of the work used to solve this problem. CH4 (g) + 2 O2 (g) yields CO2 (g) + 2 H2O (g)

OpenStudy (anonymous):

How many liters of 13.5 molar HCl stock solution are needed to make 10.5 liters of a 1.8 molar HCl solution? Show the work used to solve this problem.

OpenStudy (zepp):

2 questions in one? :|

OpenStudy (anonymous):

Sorry, I can separate them so you can get 2 medals?

OpenStudy (zepp):

No no, it's okay, lemme see.

OpenStudy (zepp):

And by 275 K you mean 275 kilojoules right?

OpenStudy (anonymous):

I actually have no idea....

OpenStudy (anonymous):

I'll just skip it

OpenStudy (anonymous):

lets just do the second one

OpenStudy (zepp):

lol ok

OpenStudy (zepp):

So from what I understand, you need to make a dilution in the second problem.

OpenStudy (anonymous):

Okay..

OpenStudy (zepp):

Um, I forgot how to dilution stuffs, I'm re-reading my Chem book, will be back with the solution shortly :$

OpenStudy (zepp):

Ok

OpenStudy (zepp):

Let's identify our variables. We have a stock solution (HCL) concentrated at 13.5M/L We need a 10.5 liters of a 1.8 molar HCl solution

OpenStudy (zepp):

We'll be using the formula \(C_1V_1=C_2V_2\) to find our solution.

OpenStudy (zepp):

So, our \(C_1\) would be \(13.5mol/L\) \(C_2=1.8mol/L\) \(V_2=10.5L\) We need to find how many liters of the stock solution of HCl we need \(C_1V_1=C_2V_2\\13.5M/L*V_1=1.8M/L * 10.5L\\13.5M/L*V_1=18.9M\\V_1=18.9M \div 13.5M/L=1.4L\)